# Weak expansion properties and a large deviation principle for coarse expanding conformal systems

### v2

@@ 3.2. Topological CXC systems
In this subsection, we recall topological CXC systems. We refer the reader to Section 2.2 in [HP09] for more details.
@@ 3.2.1. Basic assumptions and the definition
Let $f\colon\mathfrak{X}_{1}\rightarrow\mathfrak{X}_{0}$ be an FBC map of degree $\deg(f)=d\geq 2$. For each $n\in\mathbb{N}_{0}$ we define
@@
    $f|_{X}\colon X\rightarrow X$ is an FBC map of degree $d$. For convenience, we denote $f_{X}\coloneqq f|_{X}$.
Referring to Section 2.2 of [HP09], the following properties hold for $X$ and $\mathfrak{X}_{n}$, $n\in\mathbb{N}$.
@@
In this paper, following the terminology in [HP09], a preimage under $f$ of a connected set $A$ is defined as a connected component of $f^{-1}(A)$.
@@
$\mathcal{U}_{n+1}\coloneqq\{\widetilde{U}:\,\mbox{there exists }U\in\mathcal{U}_{n}\mbox{ such that }\widetilde{U}\mbox{ is a preimage of }U\}.$
The elements of $\mathcal{U}_{n}$ are connected components of $f^{-n}(U)$, where $U$ ranges over $\mathcal{U}_{0}$. We note that, by Proposition 3.6, for each $\widetilde{U}\in\mathcal{U}_{n+1}$ and $U\in\mathcal{U}_{n}$, if $\widetilde{U}$ is a preimage of $U$, then $f|_{U}\colon\widetilde{U}\rightarrow U$ is surjective, and that $f^{k}(U)\in\mathcal{U}_{n-k}$ for all $k\in\mathbb{N}_{0}$ with $k\leq n$. We can see that $\mathcal{U}_{n}$ is a finite open cover of $X$ by connected open sets in $\mathfrak{X}_{n+1}$. Note that since $\overline{\mathfrak{X}_{1}}$ is compact in $\mathfrak{X}_{0}$, each $U$ in $\bigcup_{n\in\mathbb{N}_{0}}\mathcal{U}_{n}$ is bounded.
We say that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ is a topological coarse expanding conformal system (abbr: topological CXC) with repellor $X$ if there exists a sequence of finite open cover $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ of $X$ constructed as above, such that the following axioms hold:
@@
    Expansion Axiom (abbr: [Expans]): For each finite open cover $\mathcal{V}$ of $X$ by open sets of $\mathfrak{X}_{0}$, there exists $N\in\mathbb{N}$ such that for each integer $n\geq N$ and each $U\in\mathcal{U}_{n}$, there exists $V\in\mathcal{V}$ with $U\subseteq V$.
@@
    Degree Axiom (abbr: [Deg]): The set of degrees of maps of the form $f^{k}|_{\widetilde{U}}\colon\widetilde{U}\rightarrow U$, where $U\in\mathcal{U}_{n},\ \widetilde{U}\in\mathcal{U}_{n+k}$ and $n,\,k\in\mathbb{N}$ are arbitrary, has a finite maximum, denoted by $p$.
@@ 3.2.2. Elementary properties
Below, we give some properties of topological CXC systems.
@@
Let $f\colon\mathfrak{X}_{1}\rightarrow\mathfrak{X}_{0}$ be an FBC map with $\deg(f)\geq 2$. Suppose that $f_{X}\coloneqq f|_{X}$ is an FBC map such that $\deg(f_{X})=\deg(f)$. Then for each $x\in X$, we have $\deg(f;x)=\deg(f_{X};x)$. Furthermore, $\deg(f^{n};x)=\deg(f_{X}^{n};x)$ holds for all $x\in X$ and $n\in\mathbb{N}$.
@@
By the total invariance of $X$, we have $f^{-1}(y)=f_{X}^{-1}(y)$. Since $\deg(f_{X};z)\leq\deg(f;z)$ holds for each $z\in X$, we get from (3.3) that $\deg(f;z)=\deg(f_{X};z)$ holds for each $z\in f^{-1}(y)$. Thus $\deg(f;x)=\deg(f_{X};x)$.
@@
    $\deg(f^{m}|_{\widetilde{U}})\leq\frac{p}{\deg(f^{n}|_{U})}$, where $p$ is the constant from [Deg].
@@
The following property of the post-branch set is established in [HP09, Proposition 2.4.3].
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a topological CXC system, and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ be a sequence of good open covers. Then the post-branch set $P_{f}$ is a nowhere dense subset of $X$.
@@ 3.3.1. The definition
With topological CXC systems recalled above, we give a brief introduction to metric CXC systems. We refer the reader to Section 2.5 in [HP09] for more details.
First, we recall the notion of roundness. Let $Z$ be a metric space and $A$ be a bounded proper subset of $Z$ with a nonempty interior. For each $a\in$ int$(A)$, denote
$L(A,a)$ $\coloneqq\sup\{\lvert b-a\rvert:b\in A\},$
$l(A,a)$ $\coloneqq\sup\{r\in(0,+\infty):B(a,r)\subseteq A\}.$
@@
For a number $K>0$, we say that $A$ is $K$-almost round if $\operatorname{Round}(A,a)\leq K$ for some $a\in A$. If $A$ is $K$-almost round, then there exists a real number $s>0$ such that $B(a,s)\subseteq A\subseteq\overline{B(a,Ks)}$.
Suppose that we have a topological CXC system $f\colon\mathfrak{X}_{1}\rightarrow\mathfrak{X}_{0}$ with repellor $X$ and a sequence of good open covers $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$. Assume that $\mathfrak{X}_{0}$ is endowed with a metric compatible with its topology. Then $f$ is called a metric CXC system if it satisfies the following two axioms:
@@
    Roundness Distortion Axiom (abbr: [Round]): There exist continuous increasing embeddings $\rho_{+}\colon[1,+\infty)\rightarrow[1,+\infty)$ and $\rho_{-}\colon[1,+\infty)\rightarrow[1,+\infty)$ such that for all $n,\,k\in\mathbb{N}_{0}$, $U\in\mathcal{U}_{n}$, $\widetilde{U}\in\mathcal{U}_{n+k}$, $y\in U$, and $\widetilde{y}\in\widetilde{U}$, if $f^{k}\bigl{(}\widetilde{U}\bigr{)}=U$ and $f^{k}(\widetilde{y})=y$, then
    $\operatorname{Round}(\widetilde{U},\widetilde{y})$ $<\rho_{-}(\operatorname{Round}(U,y)),$
    $\operatorname{Round}(U,y)$ $<\rho_{+}(\operatorname{Round}(\widetilde{U},\widetilde{y})).$
@@
    Diameter Distortion Axiom (abbr: [Diam]): There exist increasing homeomorphisms $\delta_{+}\colon[0,1]\rightarrow[0,1]$ and $\delta_{-}\colon[0,1]\rightarrow[0,1]$ such that for all $n_{0},\,n_{1},\,k\in\mathbb{N}_{0}$, $U\in\mathcal{U}_{n_{0}}$, $U^{\prime}\in\mathcal{U}_{n_{1}}$, $\widetilde{U}\in\mathcal{U}_{n_{0}+k}$, and $\widetilde{U}^{\prime}\in\mathcal{U}_{n_{1}+k}$, if $\widetilde{U}^{\prime}\subseteq\widetilde{U}$, $U^{\prime}\subseteq U$, $f^{k}\bigl{(}\widetilde{U}\bigr{)}=U$, and $f^{k}\bigl{(}\widetilde{U}^{\prime}\bigr{)}=U^{\prime}$, then
@@
    ${\operatorname{diam}U^{\prime}}/{\operatorname{diam}U}$ $<\delta_{+}({\operatorname{diam}\widetilde{U}^{\prime}}/{\operatorname{diam}\widetilde{U}}).$
@@ 3.3.2. Regularity of metric CXC systems
In this subsection, we recall some properties of metric CXC systems. The results are recorded below, and more details can be found in Section 2.6 of [HP09].
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a metric CXC system, and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ be a sequence of good open covers. Then there exists a constant $K>1$ and non-increasing sequences $\{c_{n}\}_{n\in\mathbb{N}_{0}}$ and $\{d_{n}\}_{n\in\mathbb{N}_{0}}$ of positive real numbers converging to 00 such that the following statements hold for all $n\in\mathbb{N}_{0}$ and $x\in X$:
@@
    $0<c_{n}\leq\inf\{\operatorname{diam}U:U\in\mathcal{U}_{n}\}\leq\sup\{\operatorname{diam}U:U\in\mathcal{U}_{n}\}\leq d_{n}$.
@@
    If $0<r<\frac{c_{n}}{2K}$, then there exist $U\in\mathcal{U}_{n}$ and $s>r$ such that
@@
    In particular, let $\delta_{n}$ be the Lebesgue number of $\mathcal{U}_{n}$, then $\delta_{n}\geq\frac{c_{n}}{2K}$.
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a metric CXC system, and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ be a sequence of good open covers. Then there exist constants $\lambda,\,\theta\in(0,1)$, and $C^{\prime}>0$ such that for all $n,\,k\in\mathbb{N}_{0}$, $U^{\prime}\in\mathcal{U}_{n+k}$, $U\in\mathcal{U}_{n}$, if $U\cap U^{\prime}\cap X\neq\emptyset$, then the following statements hold:
@@
    $\operatorname{diam}U^{\prime}/\operatorname{diam}U\leq C^{\prime}\theta^{k}$.
@@ 4. The Assumptions
We state below the assumptions under which we will develop our theory in most parts of this article. We will repeatedly refer to such assumptions later.
@@
    $f\colon\mathfrak{X}_{1}\rightarrow\mathfrak{X}_{0}$ is an FBC map of degree $\deg(f)\eqqcolon d\geq 2$ with repellor $X$, and $f|_{X}\colon X\rightarrow X$ is an FBC map of degree $d$.
@@
We note that, in this article, the terminology preimage under $f$ of a connected set $A$ refers to a connected component of $f^{-1}(A)$.
@@ 5.1. Basic concepts
which is also a cover of $Z$. Note that if $\xi$ and $\eta$ are both open covers, then $\xi\vee\eta$ is also an open cover. Define $g^{-1}(\xi)\coloneqq\{g^{-1}(A_{j}):j\in J\}$, and we denote for each $n\in\mathbb{N}$,
@@
    For each open cover $\eta$ of $Z$, there exists $j\in\mathbb{N}$ such that for each integer $i\geq j$, $\xi_{i}$ is a refinement of $\eta$.
@@
(5.1) $H(\xi|\eta)=\log\Bigl{(}\max\limits_{A\in\eta}\bigl{\{}\min\bigl{\{}\operatorname{card}\xi_{A}:\xi_{A}\subseteq\xi,\ A\subseteq\bigcup\xi_{A}\bigr{\}}\bigr{\}}\Bigr{)}.$
@@
(5.2) $h(g|\eta)\coloneqq\lim_{l\to+\infty}\lim_{n\to+\infty}\frac{1}{n}H\biggl{(}\bigvee_{i=0}^{n-1}g^{-i}(\xi_{l})\,\bigg{|}\,\bigvee_{j=0}^{n-1}g^{-j}(\eta)\biggr{)},$
@@
(5.3) $h^{*}(g)\coloneqq\lim_{m\to+\infty}\lim_{l\to+\infty}\lim_{n\to+\infty}\frac{1}{n}H\biggl{(}\bigvee_{i=0}^{n-1}g^{-i}(\xi_{l})\,\bigg{|}\,\bigvee_{j=0}^{n-1}g^{-j}(\eta_{m})\biggr{)},$

### v3

@@ 3.2. Topological CXC systems
In this subsection, we recall topological CXC systems. We refer the reader to [HP09, Section 2.2] for more details.
@@ 3.2.1. Basic assumptions and the definition
Let $f\colon\mathfrak{X}_{1}\rightarrow\mathfrak{X}_{0}$ be an FBC map of degree $\deg(f)=d\geqslant 2$. For each $n\in\mathbb{N}_{0}$ we define
@@
    $f|_{X}\colon X\rightarrow X$ is an FBC map of degree $d$.
Note that, in this paper, when $f\colon\mathfrak{X}_{1}\to\mathfrak{X}_{0}$ and the repellor $X$ are clear from the context, we write $f_{X}\coloneqq f|_{X}$ for convenience.
Referring to [HP09, Section 2.2], the following properties hold for $X$ and $\mathfrak{X}_{n}$, $n\in\mathbb{N}$.
@@
In this paper, adopting the terminology on [HP09, Page 14], we define a preimage under $f$ of a connected set $A$ as a connected component of $f^{-1}(A)$.
@@
[…]
The elements of $\mathcal{U}_{n}$ are connected components of $f^{-n}(U)$, where $U$ ranges over $\mathcal{U}_{0}$. We note that, by Proposition 3.6, for each $\widetilde{U}\in\mathcal{U}_{n+1}$ and $U\in\mathcal{U}_{n}$, if $\widetilde{U}$ is a preimage of $U$, then $f|_{U}\colon\widetilde{U}\rightarrow U$ is surjective, and that $f^{k}(U)\in\mathcal{U}_{n-k}$ for all $k\in\mathbb{N}_{0}$ with $k\leqslant n$. We can see that $\mathcal{U}_{n}$ is a finite open cover of $X$ by connected open sets in $\mathfrak{X}_{n+1}$.
We say that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ is a topological coarse expanding conformal system (abbr: topological CXC) with repellor $X$ if there exists a sequence of finite open covers $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ of $X$ constructed as above, such that the following axioms hold:
@@
    Expansion Axiom (abbr: [Expans]): For each finite open cover $\mathcal{V}$ of $X$ by open sets of $\mathfrak{X}_{0}$, there exists $N\in\mathbb{N}$ such that for each integer $n\geqslant N$ and each $U\in\mathcal{U}_{n}$, there exists $V\in\mathcal{V}$ with $U\subseteq V$.
@@
    Degree Axiom (abbr: [Deg]): The set of degrees of maps of the form $f^{k}|_{\widetilde{U}}\colon\widetilde{U}\rightarrow U$, where $U\in\mathcal{U}_{n},\ \widetilde{U}\in\mathcal{U}_{n+k}$ and $n,\,k\in\mathbb{N}$ are arbitrary, has a finite maximum denoted by $p$.
@@ 3.2.2. Elementary properties
Assume that $\mathfrak{X}_{0}$ and $\mathfrak{X}_{1}$ satisfy the assumptions in Subsection 3.2.1. Let $f\colon\mathfrak{X}_{1}\rightarrow\mathfrak{X}_{0}$ be an FBC map with $\deg(f)\geqslant 2$, and $X$ be the repellor. Suppose that $f_{X}\coloneqq f|_{X}$ is an FBC map such that $\deg(f_{X})=\deg(f)$. Then, for each $x\in X$, we have $\deg(f;x)=\deg(f_{X};x)$. Furthermore, $\deg(f^{n};x)=\deg(f_{X}^{n};x)$ holds for all $x\in X$ and $n\in\mathbb{N}$.
@@
By the total invariance of $X$, we have $f^{-1}(y)=f_{X}^{-1}(y)$. Since $\deg(f_{X};z)\leqslant\deg(f;z)$ holds for each $z\in X$, we get from (3.3) that $\deg(f;z)=\deg(f_{X};z)$ holds for each $z\in f^{-1}(y)$. Thus $\deg(f;x)=\deg(f_{X};x)$.
@@
    $\deg\bigl{(}f^{m}|_{\widetilde{U}}\bigr{)}\leqslant\frac{p}{\deg(f^{n}|_{U})}$, where $p$ is the constant from [Deg].
@@
We record the following property of the post-branch set; see [HP09, Proposition 2.4.3].
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a topological CXC system. Then the post-branch set $P_{f}$ is a nowhere dense subset of $X$.
@@ 3.3.1. The definition
With topological CXC systems recalled above, we give a brief introduction to metric CXC systems. We refer the reader to [HP09, Section 2.5] for more details.
First, we recall the notion of roundness. Let $Z$ be a metric space and $A$ be a bounded proper subset of $Z$ with a nonempty interior. For each $a\in$ int(A)(A), denote
$L(A,a)$ $\coloneqq\sup\{\lvert b-a\rvert:b\in A\}>0,$
$l(A,a)$ $\coloneqq\sup\{r:r\leqslant L(A,a)\mbox{ and }B(a,r)\subseteq A\}>0.$
@@
For a number $K>0$, we say that $A$ is $K$-almost round if $\operatorname{Round}(A,a)\leqslant K$ for some $a\in A$. If $A$ is $K$-almost round, then there exists a real number $s>0$ such that $B(a,s)\subseteq A\subseteq\overline{B(a,Ks)}$.
Suppose that we have a topological CXC system $f\colon\mathfrak{X}_{1}\rightarrow\mathfrak{X}_{0}$ with repellor $X$ and a sequence of good open covers $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$. Assume that $\mathfrak{X}_{0}$ is endowed with a metric compatible with its topology. Note that each $U$ in $\bigcup_{n\in\mathbb{N}_{0}}\mathcal{U}_{n}$ is bounded, since $\overline{\mathfrak{X}_{1}}$ is compact in $\mathfrak{X}_{0}$. Then $f$ is called a metric CXC system if it satisfies the following two axioms:
@@
[…]
    $\operatorname{Round}\bigl{(}\widetilde{U},\widetilde{y}\bigr{)}$ $<\rho_{-}(\operatorname{Round}(U,y)),$
    $\operatorname{Round}(U,y)$ $<\rho_{+}\bigl{(}\operatorname{Round}\bigl{(}\widetilde{U},\widetilde{y}\bigr{)}\bigr{)}$.
@@
    Diameter Distortion Axiom (abbr: [Diam]): There exist increasing homeomorphisms $\delta_{+}\colon[0,1]\rightarrow[0,1]$ and $\delta_{-}\colon[0,1]\rightarrow[0,1]$ such that for all $n_{0},\,n_{1},\,k\in\mathbb{N}_{0}$, $\in\mathcal{U}_{n_{0}}, U𝒰$$^{\prime}\in\mathcal{U}_{n_{1}}$, $\widetilde{U}\in\mathcal{U}_{n_{0}+k}$, and $\widetilde{U}^{\prime}\in\mathcal{U}_{n_{1}+k}$, if $\widetilde{U}^{\prime}\subseteq\widetilde{U}$, $U^{\prime}\subseteq U$, $f^{k}\bigl{(}\widetilde{U}\bigr{)}=U$, and $f^{k}\bigl{(}\widetilde{U}^{\prime}\bigr{)}=U^{\prime}$, then
@@
    ${\operatorname{diam}U^{\prime}}/{\operatorname{diam}U}$ $<\delta_{+}\bigl{(}{\operatorname{diam}\widetilde{U}^{\prime}}/{\operatorname{diam}\widetilde{U}}\bigr{)}$.
@@ 3.3.2. Regularity of metric CXC systems
In this subsection, we recall some properties of metric CXC systems. The results are recorded below, and more details can be found in [HP09, Section 2.6].
The following proposition summarizes the roundness properties given by [HP09, Propositions 2.6.2 and 2.6.3].
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a metric CXC system, and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ be a sequence of good open covers. Then there exists a constant $K>1$ and non-increasing sequences $\{c_{n}\}_{n\in\mathbb{N}_{0}}$ and $\{d_{n}\}_{n\in\mathbb{N}_{0}}$ of positive real numbers converging to 0 such that the following statements hold for all $n\in\mathbb{N}_{0}$ and $x\in X$:
@@
    $0<c_{n}\leqslant\inf\{\operatorname{diam}U:U\in\mathcal{U}_{n}\}\leqslant\sup\{\operatorname{diam}U:U\in\mathcal{U}_{n}\}\leqslant d_{n}$.
@@
    If $0<r<\frac{1}{2K}\min\{\operatorname{diam}U:U\in\mathcal{U}_{n}\}$, then there exist $U\in\mathcal{U}_{n}$ and $s>r$ such that
@@
    In particular, let $\delta_{n}$ be the Lebesgue number of $\mathcal{U}_{n}$, then $\delta_{n}\geqslant\frac{c_{n}}{2K}$.
The following proposition summarizes the local comparability and exponential contraction properties given by [HP09, Propositions 2.6.4 and 2.6.5].
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a metric CXC system, and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ be a sequence of good open covers. Then there exist constants $\lambda,\,\theta\in(0,1)$, and $C^{\prime}>0$ such that for all $n,\,k\in\mathbb{N}_{0}$, $^{\prime}\in\mathcal{U}_{n+k}, U𝒰$$\in\mathcal{U}_{n}$, if $U\cap U^{\prime}\cap X\neq\emptyset$, then the following statements hold:
@@
    $\operatorname{diam}U^{\prime}/\operatorname{diam}U\leqslant C^{\prime}\theta^{k}$.
@@ 4. The Assumptions
We state below the assumptions under which we will develop our theory in most parts of this paper. We will repeatedly refer to such assumptions later.
@@
    $f\colon\mathfrak{X}_{1}\rightarrow\mathfrak{X}_{0}$ is an FBC map of degree $\deg(f)\eqqcolon d\geqslant 2$ with repellor $X$, and $f|_{X}\colon X\rightarrow X$ is an FBC map of degree $d$.
@@
We fix some notations for this paper. In this paper, when $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the assumptions above, we write $f_{X}\coloneqq f|_{X}$, and
(4.1) $\mathcal{W}_{n}\coloneqq\{U\cap X:U\in\mathcal{U}_{n}\},$
for each $n\in\mathbb{N}_{0}$. We note that, in this paper, the terminology preimage under $f$ of a connected set $A$ refers to a connected component of $f^{-1}(A)$.
@@ 5.1. Basic concepts
which is also a cover of $Z$. Note that if $\xi$ and $\eta$ are both open covers, then $\xi\vee\eta$ is also an open cover. Define $g^{-1}(\xi)\coloneqq\bigl{\{}g^{-1}(A_{j}):j\in J\bigr{\}}$, and we denote for each $n\in\mathbb{N}$,
@@
    For each open cover $\eta$ of $Z$, there exists $j\in\mathbb{N}$ such that for each integer $i\geqslant j$, $\xi_{i}$ is a refinement of $\eta$.
@@
(5.1) $H(\xi|\eta)=\log\Bigl{(}\max\limits_{A\in\eta}\bigl{\{}\min\bigl{\{}\operatorname{card}\xi_{A}:\xi_{A}\subseteq\xi,\ A\subseteq\bigcup\xi_{A}\bigr{\}}\bigr{\}}\Bigr{)}$
and
(5.2) $h(g,\xi|\eta)\coloneqq\lim_{n\to+\infty}\frac{1}{n}H\biggl{(}\bigvee_{i=0}^{n-1}g^{-i}(\xi)\,\bigg{|}\,\bigvee_{j=0}^{n-1}g^{-j}(\eta)\biggr{)}.$
@@
(5.3) $h(g|\eta)\coloneqq\lim_{l\to+\infty}h(g,\xi_{l}|\eta),$
@@
(5.4) $h^{*}(g)\coloneqq\lim_{m\to+\infty}\lim_{l\to+\infty}h(g,\xi_{l}|\eta_{m}),$
@@
Definition 5.3 (Forward expansive).
A continuous map $g\colon Z\rightarrow Z$ on compact metric space $Z$ is forward expansive if there exists $\delta>0$ such that for all distinct points $x,\,y\in Z$, there exists $n\in\mathbb{N}_{0}$ for which $d(f^{n}(x),f^{n}(y))\geqslant\delta$.
@@
Remark 5.5.
It is shown in [Mi76] that a continuous map $g\colon Z\to Z$ on compact metric space $(Z,d)$ is $h$-expansive if and only if
$\sup_{x\in Z}\biggl{\{}\lim_{\delta\to 0}\limsup_{n\to+\infty}\frac{1}{n}\log r_{n}(g,\Phi_{\epsilon}(x),\delta)\biggr{\}}=0$
for some $\epsilon>0$. Here $\Phi_{\epsilon}(x)$ is defined as
$\Phi_{\epsilon}(x)\coloneqq\bigl{\{}y\in Z:d(g^{n}(x),g^{n}(y))\leqslant\epsilon\mbox{ for each }n\in\mathbb{N}_{0}\bigr{\}},$
and $r_{n}(g,\Phi_{\epsilon}(x),\delta)$ stands for the smallest cardinality of a set that $(n,\delta)$-spans $\Phi_{\epsilon}(x)$. If $g$ is forward expansive, then there is $\epsilon>0$ such that $\Phi_{\epsilon}(x)=\{x\}$ for each $x\in Z$, which implies that $g$ is $h$-expansive.
@@
Remark.
It is shown in [Mi76] that a continuous map $g\colon Z\to Z$ on compact metric space $Z$ is asymptotic $h$-expansive whenever $g$ is $h$-expansive.